The centers of a 7.0 kg lead ball and a 50 g lead ball are separated by 12cm. What is the ratio of this gravitational force to the weight of the 50 g ball?
Answer #1

Calculate the gravitational force = G*m1*m2/r^2

F = 6.67*(10^-11)*7*0.05/(0.12*0.12)

= 1.62*(10^-9) N

Calculate weight, W = 0.05 *9.81

ratio of gravitational force to weight.

= (1.62*(10^-9))/ 0.05*9.8

= 3.306 *10^-9 answer.

Please rate positive if it is helpful